what is the missing term of the following solution?..-34,_,77
Answer by swincher4391(1107)

I'm going to assume that your sequence is just adding by some common difference d. If that's the case, then take the midpoint between -34 and 77. Then you get -34 +77 / 2 = 43/2 = 21.5. Without more information, this is all I can gather.